# Decentralized Stochastic Subgradient-based Methods for Nonsmooth Nonconvex Optimization

### v2

@@ 1.4 Organizations
The rest of this paper is organized as follows. In section 2, we present some basic notations and definitions. In section 3, we describe our proposed framework (DSM) in detail and establish the global convergence by linking it with the dynamics of its corresponding differential inclusion. In Section 4, we demonstrate that framework (DSM) covers the major existing decentralized stochastic subgradient-based methods. Numerical experiments and results of analyzed decentralized subgradient-based methods are described in Section 5. In the last section, we draw conclusions and discuss possible future research directions.
@@ 2.1 Notations
For any positive sequence $\left\{\eta_{k}\right\}$, let $\lambda_{\eta}(0):=0,\lambda_{\eta}(i):=\sum_{k=0}^{i-1}\eta_{k}$, and $\Lambda_{\eta}(t):=\sup\left\{k\in\mathbb{N}:t\geq\lambda_{\eta}(k)\right\}.$ More explicitly, $\Lambda_{\eta}(t)=p$, if $\lambda_{\eta}(p)\leq t<\lambda_{\eta}(p+1)$. Furthermore, we define the set-valued mapping sign : $\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ as follows:
@@
Then it is easy to verify that $\operatorname{sign}({\bm{x}})$ is the Clarke subdifferential of the nonsmooth functions ${\bm{x}}\mapsto\|{\bm{x}}\|_{1}$.
In addition, we denote $\mathcal{N}_{i}$ as the set of neighbor agents of agent $i$ together with $\{i\}$, $\mathbb{R}_{+}$ as the set of all nonnegative real numbers. For any $N>0,[N]:=\{1,\ldots,N\}$. And let ${\bm{1}}_{d}$ represent all-ones vector.
Furthermore, we denote $(\Omega,\mathcal{F},\mathbb{P})$ as the probability space. We say $\left\{\mathcal{F}_{k}\right\}_{k\in\mathbb{N}}$ is a filtration if $\left\{\mathcal{F}_{k}\right\}$ is a collection of $\sigma$-algebras that satisfies $\mathcal{F}_{0}\subseteq\mathcal{F}_{1}\subseteq\cdots\subseteq\mathcal{F}_{\infty}\subseteq\mathcal{F}$. A stochastic series $\{\xi_{k}\}$ is called a martingale with respect to a filtration $\{\mathcal{F}_{k}\}$, if $\{\xi_{k}\}$ is adapted to the filtration $\{\mathcal{F}_{k}\}$ and $\mathbb{E}[\xi_{k+1}|\mathcal{F}_{k}]=\xi_{k}$, $\forall k\in\mathbb{N}$. Moreover, A stochastic series $\{\xi_{k}\}$ is referred to as a martingale difference sequence, if $\{\xi_{k}\}$ is adapted to the filtration $\{\mathcal{F}_{k}\}$ and $\mathbb{E}[\xi_{k+1}|\mathcal{F}_{k}]=0$ holds for all $k\in\mathbb{N}$.
@@ 2.2 Mixing matrix
In this subsection, we introduce the concept of mixing matrix ${\bm{W}}=[{\bm{W}}(i,j)]\in\mathbb{R}^{d\times d}$, which determines the structure of the communication network and plays an important role in averaging the local information of neighbor agents. Generally, we assume the mixing matrix ${\bm{W}}$ satisfies the following properties, which are standard in the literature.
@@
[44, Section 1] Given a connected graph $\mathtt{G}=(\mathtt{V},\mathtt{E})$, we say ${\bm{W}}\in\mathbb{R}^{d\times d}$ is a mixing matrix of $\mathtt{G}$, if it satisfies the following conditions.
@@
Invoking [49, Perron-Frobenius Theorem], we derive a direct corollary, which illustrates the fundamental spectral property of mixing matrix ${\bm{W}}$.
@@ 2.3 Nonsmooth analysis and conservative field
In this subsection, we present a brief introduction to some basic concepts in nonsmooth analysis and the conservative field [28].
A set-valued mapping $D:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is a mapping from $\mathbb{R}^{n}$ to set of subsets of $\mathbb{R}^{m}$. The graph of $D$ is given by $\operatorname{graph}D=\{({\bm{x}},{\bm{z}})\mid{\bm{x}}\in\mathbb{R}^{n},{\bm{z}}\in D({\bm{x}})\}$. The set-valued mapping is said to be graph-closed (or has closed graph), if $\operatorname{graph}D$ is a closed subset of $\mathbb{R}^{n}\times\mathbb{R}^{m}$. $D$ is said to be locally bounded, if for any ${\bm{x}}\in\mathbb{R}^{n}$, there exists a neighborhood $U_{{\bm{x}}}$ of ${\bm{x}}$ such that $\sup_{{\bm{z}}\in D({\bm{y}}),{\bm{y}}\in U_{{\bm{x}}}}\|{\bm{z}}\|<+\infty$. In addition, $D$ is convex-valued, if $D({\bm{x}})$ is a convex subset of $\mathbb{R}^{n}$ for any ${\bm{x}}\in\mathbb{R}^{n}$.
Moreover, A curve $\gamma:\mathbb{R}_{+}\rightarrow\mathbb{R}^{n}$ is absolutely continuous if $\gamma$ is differentiable almost everywhere, and $\gamma(t)=\gamma(0)+\int_{0}^{t}\gamma^{\prime}(u)\mathrm{d}u$ holds for all $t\in\mathbb{R}_{+}$.
For any locally Lipschitz continuous function $f:\mathbb{R}^{n}\to\mathbb{R}$, Rademacher’s theorem states that a locally Lipschitz continuous function $f$ is differentiable almost everywhere. We denote the subset in $\mathbb{R}^{n}$ on which $f$ is differentiable by $\Omega_{f}$. According to [50, Theorem 2.5.1], we have
$\partial{f({\bm{x}})}=\operatorname{conv}\left\{\lim_{i\to+\infty}\nabla f\left({\bm{x}}_{i}\right):{\bm{x}}_{i}\rightarrow{\bm{x}},~{}{\bm{x}}_{i}\in\Omega_{f}\right\}.$
Then it is easy to verify that $\partial f$ is a convex-valued graph-closed locally bounded set-valued mapping from $\mathbb{R}^{n}$ to the subsets of $\mathbb{R}^{n}$.
Now we present a brief introduction to the conservative field.
@@
Let $D:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ be a nonempty set-valued mapping. We call $D$ a conservative field if it is compact-valued and graph closed, and for any absolutely continuous loop $\gamma:[0,1]\rightarrow\mathbb{R}^{n}$ satisfying $\gamma(0)=\gamma(1)$, it holds that
$\int_{0}^{1}\max_{{\bm{v}}\in D(\gamma(t))}\langle\dot{\gamma}(t),{\bm{v}}\rangle\mathrm{d}t=0.$ (2.1)
It is remarkable to note that any conservative field is locally bounded [28, Remark 3]. We now introduce the definition of the potential function corresponding to the conservative field.
@@
Let $D$ be a conservative field, we say a function $f$ is a potential for $D$, provided that $f$ can be formulated as
@@
for any absolutely continuous curve $\gamma$ with $\gamma(0)={\bm{x}}_{0}$ and $\gamma(t)={\bm{x}}$. We also say that $D$ is a conservative field for $f$, denoted as $D_{f}$. Such functions $f$ are called path-differentiable functions.
As illustrated in [28, Corollary 1], for any given path-differentiable function $f$, $\partial f$ is a conservative field for $f$. Therefore, the concept of the conservative field can be regarded as an extension of the Clarke subdifferential. More importantly, when the value of conservative field $D_{f}({\bm{x}})$ is a convex subset of $\mathbb{R}^{m}$, ${\bm{0}}\in\partial f({\bm{x}})$ implies that $0\in D_{f}({\bm{x}})$.
@@
Let $f$ in (1.1) be a path-differentiable function that admits $\mathcal{D}_{f}$ as its convex-valued conservative field, then we say ${\bm{X}}\in\mathbb{R}^{n\times d}$ is a $\mathcal{D}_{f}$-critical point of (DOP), if ${\bm{X}}=\frac{1}{d}{\bm{X}}{\bm{1}_{d}}{\bm{1}}_{d}^{\top}$ and ${\bm{0}}\in\mathcal{D}_{f}(\frac{1}{d}{\bm{X}}{\bm{1}_{d}})$.
2.4 Differential inclusion and stochastic approximation
In this subsection, we introduce fundamental concepts and theories associated with differential inclusion, which is crucial for establishing the convergence properties of stochastic approximation. We begin with the concept of differential inclusion.
@@ 2.4 Differential inclusion and stochastic approximation
We now present the definitions of the Lyapunov function and perturbed solution.
@@
For any set-valued mapping $D:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ and any $\delta\geq 0$, we denote $D^{\delta}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ as $D^{\delta}({\bm{x}})=\cup_{{\bm{y}}\in\mathbb{B}_{\delta}({\bm{x}})}(D({\bm{y}})+\mathbb{B}(\delta,{\bm{0}}))$.
@@
  • For above locally integrable $w$, there exists $\delta:\mathbb{R}_{+}\rightarrow\mathbb{R}$ such that $\lim_{t\rightarrow+\infty}\delta(t)=0$ and $\dot{\gamma}(t)-w(t)\in D^{\delta(t)}(\gamma(t))$.
Then, we consider a discrete iterative sequence corresponding to the differential inclusion (2.2),
${\bm{x}}_{k+1}={\bm{x}}_{k}+\eta_{k}({\bm{y}}_{k}+\xi_{k+1}),$ (2.3)
where $\{\eta_{k}\}$ is a diminishing positive sequence of step-sizes, ${\bm{y}}_{k}$ is an approximate estimate of $D$ at ${\bm{x}}_{k}$, and $\xi_{k+1}$ is a random noise when evaluating ${\bm{y}}_{k}$. A continuous time interpolated process $u:\mathbb{R}_{+}\rightarrow\mathbb{R}^{n}$ can be induced by (2.3),
$u(\lambda_{\eta}(k)+s)={\bm{x}}_{k}+\frac{{\bm{x}}_{k+1}-{\bm{x}}_{k}}{\eta_{k}}s,\quad s\in[0,\eta_{k}).$ (2.4)
Here, $\lambda_{\eta}(0):=0$ and $\lambda_{\eta}(n):=\sum_{i=0}^{n-1}\eta_{i},n\geq 1$.
The following theorem summarizes the results in [30], which illustrates the convergence of $\{{\bm{x}}_{k}\}$ generated by (2.3). Therefore, we omit the proof of Theorem 2.10 for simplicity.
Theorem 2.10.
Let $D:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ be a locally bounded set-valued mapping that is nonempty compact convex valued with closed graph. For any sequence $\{{\bm{x}}_{k}\}$, suppose there exists a continuous function $\psi:\mathbb{R}^{n}\to\mathbb{R}$ and a closed subset $\mathcal{A}$ of $\mathbb{R}^{n}$ such that
@@
    $\psi$ is bounded from below, and the set $\{\psi({\bm{x}})\,:\,{\bm{x}}\in\mathcal{A}\}$ has empty interior in $\mathbb{R}$.
@@
    $\psi$ is a Lyapunov function for the differential inclusion (2.2) that admits $\mathcal{A}$ as its stable set.
@@
    The interpolated process of $\{{\bm{x}}_{k}\}$ is a perturbed solution of (2.2).
Then any cluster point of $\{{\bm{x}}_{k}\}$ lies in $\mathcal{A}$, and the sequence $\{\psi({\bm{x}}_{k})\}$ converges.
Similar results under slightly different conditions can be found in [31, 29, 32], while some recent works [51, 52] focus on analyzing the convergence of (2.3) under relaxed conditions. Interested readers could refer to those works for details.
In this section, we aim to establish the global convergence of the framework (DSM). Section 3.1 exhibits the fundamental assumptions in our analysis, and in Section 3.2 we present the convergence properties of the framework (DSM).
3.1 Basic assumptions
In this subsection, we introduce the basic assumptions of our theoretical analysis. To begin with, we make some mild assumptions on our framework (DSM).
@@ 3.1 Basic assumptions
    The sequences $\{{\bm{Z}}_{k}\}$ and $\{{\bm{H}}_{k}\}$ are bounded.
@@
    The sequence of step-sizes $\{\eta_{k}\}$ is positive and diminishing, which satisfies $\sum_{k=0}^{+\infty}\eta_{k}=+\infty$. Moreover, for any $T>0$, it holds that
    $\lim_{i\to+\infty}\sup_{i\leq j\leq\Lambda_{\eta}(\lambda_{\eta}(i)+T)}\left\|\sum_{k=i}^{j}\eta_{k}\Xi_{k+1}\right\|=0.$
  3. 3.
    There exists a positive diminishing sequence $\{\delta_{k}\}$ and a locally bounded graph-closed set-valued mapping $\Phi_{i}$, such that for each $i\in[d]$, it holds that
Remark 3.2.
Assumption 3.1(1) assumes the uniform boundedness of the sequence $\{{\bm{Z}}_{k}\}$, which is commonly employed in existing works [28, 29, 36, 39]. Assumption 3.1(2) illustrates that the noises in the framework (DSM) can be controlled by the step-sizes, which is identical to the condition on the noises in [30]. Furthermore, Assumption 3.1(3) illustrates that the averaged update directions of all the agents $\{\frac{1}{d}{\bm{H}}_{k}{\bm{1}_{d}}\}$ approximate the evaluation of $\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}({\bm{z}}_{i,k})$.
For conciseness, we define the set-valued mapping $\Phi:\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ as $\Phi({\bm{z}}):=\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}({\bm{z}})\right)$. Then we present the following assumptions on the differential inclusion
$\frac{\mathrm{d}{\bm{z}}}{\mathrm{d}t}\in-\Phi({\bm{z}}),$ (3.2)
which corresponds to the discrete update scheme in (DSM).

### v3

@@ 1.5 Organizations
The rest of this paper is organized as follows. In Section 2, we present some basic notations and definitions. In Section 3, we stipulate some assumptions for our proposed framework (DSM), and establish convergence results under different noise settings by linking it with the dynamics of its corresponding differential inclusion. In Section 4, we demonstrate that framework (DSM) covers major existing decentralized stochastic subgradient-based methods, which inherit the convergence properties from (DSM) and avoid infinitely many spurious critical points. Numerical experiments regarding analyzed decentralized subgradient-based methods are presented in Section 5. In the last section, we draw conclusions and discuss possible future research directions.
@@ 2.1 Notations
For any positive sequence $\left\{\eta_{k}\right\}$, let $\lambda_{\eta}(0):=0,\lambda_{\eta}(i):=\sum_{k=0}^{i-1}\eta_{k}$, and $\Lambda_{\eta}(t):=\sup\left\{k\in\mathbb{N}:\right.$ $\left.t\geq\lambda_{\eta}(k)\right\}.$ More explicitly, $\Lambda_{\eta}(t)=p$, if $\lambda_{\eta}(p)\leq t<\lambda_{\eta}(p+1)$. We define the set-valued mapping sign : $\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ as follows:
@@
It is easy to verify that $\operatorname{sign}({\bm{x}})$ is the Clarke subdifferential of the nonsmooth functions ${\bm{x}}\mapsto\|{\bm{x}}\|_{1}$.
Moreover, we denote $\mathcal{N}_{i}$ as the set of neighbor agents of agent $i$ together with $\{i\}$, $\mathbb{R}_{+}$ as the set of all nonnegative real numbers. For any $N>0,[N]:=\{1,\ldots,N\}$. And let ${\bm{1}}_{d}\in\mathbb{R}^{d}$ represent a vector of all $1$’s, ${\bm{e}}_{i}\in\mathbb{R}^{d}$ represent $[0,\ldots,1,\ldots,0]^{\top}$, where $1$ is the $i$-th component. For two integers $i$ and $j$, $i\land j$ represents $\min\{i,j\}$.
We denote $(\Omega,\mathcal{F},\mathbb{P})$ as the probability space. We use $\sigma({\bm{X}})$ to denote the sigma-algebra generated by the random variable ${\bm{X}}$. We say $\left\{\mathcal{F}_{k}\right\}_{k\in\mathbb{N}}$ is a filtration if $\left\{\mathcal{F}_{k}\right\}$ is a collection of $\sigma$-algebras that satisfies $\mathcal{F}_{0}\subseteq\mathcal{F}_{1}\subseteq\cdots\subseteq\mathcal{F}_{\infty}\subseteq\mathcal{F}$. A stochastic series $\{\xi_{k}\}$ is called a martingale with respect to a filtration $\{\mathcal{F}_{k}\}$, if $\{\xi_{k}\}$ is adapted to the filtration $\{\mathcal{F}_{k}\}$ and $\mathbb{E}[\xi_{k+1}|\mathcal{F}_{k}]=\xi_{k}$, $\forall k\in\mathbb{N}$. Moreover, A stochastic series $\{\xi_{k}\}$ is referred to as a martingale difference sequence, if $\{\xi_{k}\}$ is adapted to the filtration $\{\mathcal{F}_{k}\}$ and $\mathbb{E}[\xi_{k+1}|\mathcal{F}_{k}]=0$ holds for all $k\in\mathbb{N}$.
In addition, we define the summation function $f$ of (DOP) as
$f({\bm{x}}):=\frac{1}{d}\sum_{i=1}^{d}f_{i}({\bm{x}}).$ (2.1)
@@ 2.2 Mixing matrix
Mixing matrix ${\bm{W}}=[{\bm{W}}(i,j)]\in\mathbb{R}^{d\times d}$ determines the structure of the communication network and plays an important role in averaging the local information of neighbor agents. Generally, we assume the mixing matrix ${\bm{W}}$ satisfies the following properties, which are standard in the literature [52, 53, 9].
@@
Given a connected graph $\mathtt{G}=(\mathtt{V},\mathtt{E})$, we say ${\bm{W}}\in\mathbb{R}^{d\times d}$ is a mixing matrix of $\mathtt{G}$, if it satisfies the following conditions.
@@
Based on [57, Perron-Frobenius Theorem], we derive a direct corollary, which illustrates the fundamental spectral property of mixing matrix ${\bm{W}}$.
@@ 2.3 Nonsmooth analysis and conservative field
A set-valued mapping $D:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ is a mapping from $\mathbb{R}^{n}$ to set of subsets of $\mathbb{R}^{n}$. The graph of $D$ is given by $\operatorname{graph}D=\{({\bm{x}},{\bm{z}})\mid{\bm{x}}\in\mathbb{R}^{n},{\bm{z}}\in D({\bm{x}})\}$. The set-valued mapping is said to be graph-closed (or has closed graph), if $\operatorname{graph}D$ is a closed subset of $\mathbb{R}^{n}\times\mathbb{R}^{n}$. $D$ is said to be locally bounded, if for any ${\bm{x}}\in\mathbb{R}^{n}$, there exists a neighborhood $U_{{\bm{x}}}$ of ${\bm{x}}$ such that $\sup_{{\bm{z}}\in D({\bm{y}}),{\bm{y}}\in U_{{\bm{x}}}}\|{\bm{z}}\|<+\infty$. In addition, $D$ is convex-valued, if $D({\bm{x}})$ is a convex subset of $\mathbb{R}^{n}$ for any ${\bm{x}}\in\mathbb{R}^{n}$.
Definition 2.3.
[33] For any given locally Lipschitz continuous function $f:\mathbb{R}^{n}\to\mathbb{R}$, and any ${\bm{x}}\in\mathbb{R}^{n}$, the generalized directional derivative of $f$ at ${\bm{x}}$ along the direction ${\bm{d}}\in\mathbb{R}^{n}$, is defined as
$f^{\circ}({\bm{x}};{\bm{d}}):=\limsup_{{\bm{y}}\to{\bm{x}},t\downarrow 0}\frac{f({\bm{y}}+t{\bm{d}})-f({\bm{y}})}{t}.$
The Clarke subdifferential of $f$ at ${\bm{x}}\in\mathbb{R}^{n}$, denoted by $\partial f({\bm{x}})$, is defined as
$\partial f({\bm{x}}):=\left\{{\bm{u}}\in\mathbb{R}^{n}:f^{\circ}({\bm{x}};{\bm{d}})\geq\langle{\bm{u}},{\bm{d}}\rangle,\forall{\bm{d}}\in\mathbb{R}^{n}\right\}.$
Notice that $\partial f$ is a convex-valued graph-closed locally bounded set-valued mapping from $\mathbb{R}^{n}$ to the subsets of $\mathbb{R}^{n}$. Based upon the concept of generalized directional derivative, we present the definition of Clarke regular functions.
Definition 2.4.
[33] We say that $f$ is Clarke regular at ${\bm{x}}\in\mathbb{R}^{n}$, if for every direction ${\bm{d}}\in\mathbb{R}^{n}$, the one-sided directional derivative
$f^{*}({\bm{x}};{\bm{d}}):=\lim_{t\downarrow 0}\frac{f({\bm{x}}+t{\bm{d}})-f({\bm{x}})}{t}$
exists, and $f^{*}({\bm{x}};{\bm{d}})=f^{\circ}({\bm{x}};{\bm{d}}).$
Next, we present a brief introduction to the conservative field, which characterizes the output of AD algorithms in differentiating nonsmooth neural networks.
@@
Let $D:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ be a nonempty set-valued mapping. We call $D$ a conservative field if it is compact-valued and graph-closed, and for any absolutely continuous loop $\gamma:[0,1]\rightarrow\mathbb{R}^{n}$ satisfying $\gamma(0)=\gamma(1)$, it holds that
$\int_{0}^{1}\max_{{\bm{v}}\in D(\gamma(t))}\langle\dot{\gamma}(t),{\bm{v}}\rangle\mathrm{d}t=0.$
It is remarkable to note that any conservative field is locally bounded [36, Remark 3]. We now introduce the definition of the path-differentiable function corresponding to the conservative field.
@@
Let $D$ be a conservative field, we say a function $f$ is a potential for $D$ or $f$ is path-differentiable, provided that $f$ can be formulated as
@@
for any absolutely continuous curve $\gamma$ with $\gamma(0)={\bm{x}}_{0}$ and $\gamma(t)={\bm{x}}$. We also say that $D$ is a conservative field for $f$, denoted as $D_{f}$.
As illustrated in [36, Corollary 1], for any given path-differentiable function $f$, $\partial f$ is a conservative field for $f$. Therefore, the concept of the conservative field can be regarded as an extension of the Clarke subdifferential. More importantly, when conservative field $D_{f}({\bm{x}})$ is convex-valued, ${\bm{0}}\in\partial f({\bm{x}})$ implies that $0\in D_{f}({\bm{x}})$.
In fact, the class of path-differentiable functions is general enough to cover the most objective functions in real-world problems. The well-known Clarke regular functions mentioned in [33] and semi-algebraic functions mentioned in [58] are all path-differentiable functions. In [37, 36], the authors identify definable functions (whose graphs are definable in an o-minimal structure) as an important class of path-differentiable functions. Actually, most activation functions and loss functions in deep neural networks are definable functions, including sigmoid, softplus, ReLU, leaky ReLU, hinge loss, etc.
Furthermore, it is important to highlight that definability is preserved under finite summation and composition [37]. Consequently, for any neural network constructed by definable blocks, its loss function is definable, thus it is a path-differentiable function. Additionally, it is worth noting that the Clarke subdifferential of definable functions is itself definable [36]. Hence, in the case of a neural network constructed by definable blocks, the conservative field corresponding to the AD algorithms is also definable. The following proposition demonstrates that the definability of both $f$ and its conservative field $D_{f}$ leads to the nonsmooth Morse–Sard property [59].
Proposition 2.7 (Theorem 5 in [36]).
Let $f$ be a path-differentiable function that admits $D_{f}$ as its conservative field. Suppose both $f$ and $D_{f}$ are definable over $\mathbb{R}^{n}$, then the set $\left\{{f({\bm{x}})}:{\bm{0}}\in D_{f}({\bm{x}})\right\}$ is finite.
@@
Let $f$ in (2.1) be a path-differentiable function that admits $D_{f}$ as its convex-valued conservative field, then we say ${\bm{X}}\in\mathbb{R}^{n\times d}$ is a $D_{f}$-critical point of (DOP), if ${\bm{X}}=\frac{1}{d}{\bm{X}}{\bm{1}_{d}}{\bm{1}}_{d}^{\top}$ and ${\bm{0}}\in D_{f}(\frac{1}{d}{\bm{X}}{\bm{1}_{d}})$. Furthermore, we say ${\bm{X}}\in\mathbb{R}^{n\times d}$ is a Clarke (or $\partial f$)-critical point of (DOP), if ${\bm{X}}=\frac{1}{d}{\bm{X}}{\bm{1}_{d}}{\bm{1}}_{d}^{\top}$ and ${\bm{0}}\in\partial f(\frac{1}{d}{\bm{X}}{\bm{1}_{d}})$.
2.4 Differential inclusion and stochastic subgradient methods
In this part, we introduce fundamental concepts and theories associated with differential inclusion, which is crucial for establishing the convergence properties for stochastic subgradient methods.
@@ 2.4 Differential inclusion and stochastic subgradient methods
For any set-valued mapping $D:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ and any $\delta\geq 0$, we denote $D^{\delta}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ as $D^{\delta}({\bm{x}})=\cup_{{\bm{y}}\in\mathbb{B}({\bm{x}},\delta)}(D({\bm{y}})+\mathbb{B}({\bm{0}},\delta))$.
@@
  • For above locally integrable $w$, there exists $\delta:\mathbb{R}_{+}\rightarrow\mathbb{R}$ such that $\lim_{t\rightarrow\infty}\delta(t)=0$ and $\dot{\gamma}(t)-w(t)\in D^{\delta(t)}(\gamma(t))$.
Now we consider an iterative sequence corresponding to the differential inclusion (2.2),
${\bm{x}}_{k+1}={\bm{x}}_{k}+\eta_{k}(D^{\delta_{k}}({\bm{x}}_{k})+\xi_{k+1}),$ (2.3)
where $\{\eta_{k}\}$ is a non-summable positive sequence of step-sizes, $\{\delta_{k}\}$ is a nonnegative sequence, and $\xi_{k+1}$ is a random noise when evaluating $D({\bm{x}}_{k})$. A continuous-time interpolated process $u:\mathbb{R}_{+}\rightarrow\mathbb{R}^{n}$ induced by (2.3) is defined as
$u(\lambda_{\eta}(k)+s)={\bm{x}}_{k}+\frac{{\bm{x}}_{k+1}-{\bm{x}}_{k}}{\eta_{k}}s,\quad s\in[0,\eta_{k}).$
Here, $\lambda_{\eta}(0):=0$ and $\lambda_{\eta}(k):=\sum_{i=0}^{k-1}\eta_{i},k\geq 1$.
In what follows, we summarize and present several lemmas, which respectively describe the conditions under which the interpolated process of the sequence $\{{\bm{x}}_{k}\}$ is a perturbed solution of (2.2), how the sequence $\{{\bm{x}}_{k}\}$ remains uniformly bounded, and a result on global convergence.
Lemma 2.13 (Extension of Proposition 1.3 in [45]).
Let $D:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ be a locally bounded, graph-closed and convex-valued set-valued mapping. Assume that the following hold:
@@
    For any $T>0$,
    $\lim_{s\to\infty}\sup_{s\leq i\leq\Lambda_{\eta}(\lambda_{\eta}(s)+T)}\left\|\sum_{k=s}^{i}\eta_{k}\xi_{k+1}\right\|=0,$
    where $\Lambda_{\eta}(t):=\sup\{k:\lambda_{\eta}(k)\leq t\}$.
@@
    $\{\delta_{k}\}$ is diminishing, i.e. $\lim_{k\to\infty}\delta_{k}=0$.
@@
    $\sup_{k\in\mathbb{N}}\|{\bm{x}}_{k}\|<+\infty$.
Then the interpolated process of sequence $\{{\bm{x}}_{k}\}$ is a perturbed solution for (2.3).
Before exhibiting the results on uniform boundedness and convergence of (2.3), we provide an assumption about differential inclusion (2.2), which is commonly used in the analysis of ODE-based approaches [46, 47, 48].
@@
    There exists a locally Lipschitz continuous Lyapunov function $\psi:\mathbb{R}^{n}\to\mathbb{R}$ for the differential inclusion (2.2) with stable set $\mathcal{A}$.
@@
    The set $\{\psi({\bm{x}}):{\bm{x}}\in\mathcal{A}\}$ is a finite subset of $\mathbb{R}$.
Lemma 2.15 (Theorem 3.6 in [43]).
Let $D:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ be a locally bounded, graph-closed and convex-valued set-valued mapping, and $\mathcal{X}_{0}$ be any compact subset of $\mathbb{R}^{n}$. Suppose Assumption 2.14 holds and Lyapunov function $\psi$ is coercive.
Then for any given $r>\max\{0,4\sup_{x\in\mathcal{X}_{0}\cup\mathcal{A}}\psi(x)\}$, there exist $\alpha>0$, $T>0$ such that whenever $\sup_{k\geq 0}\eta_{k}\leq\alpha$, $\sup_{k\geq 0}\delta_{k}\leq\alpha$, and $\sup_{s\leq i\leq\Lambda_{\eta}(\lambda_{\eta}(s)+T)}\|{\sum_{k=s}^{i}\eta_{k}\xi_{k+1}}\|\leq\alpha,\forall s\geq 0$, the sequence $\{x_{k}\}$ generated by (2.3) with $x_{0}\in\mathcal{X}_{0}$ is restricted in $\mathcal{L}_{r}:=\{{\bm{x}}\in\mathbb{R}^{n}:\psi({\bm{x}})\leq r\}$.
Lemma 2.16 (Summary of Theorem 3.6, Proposition 3.27 in [45]).
Let $D:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ be a locally bounded, graph-closed, and convex-valued set-valued mapping. Suppose Assumption 2.14 holds, and $\{{\bm{x}}_{k}\}$ is generated by (2.3) with any ${\bm{x}}_{0}\in\mathbb{R}^{n}$. If the interpolated process of $\{{\bm{x}}_{k}\}$ is a perturbed solution of (2.2), then any cluster point of $\{{\bm{x}}_{k}\}$ generated by (2.3) lies in $\mathcal{A}$, and the sequence $\{\psi({\bm{x}}_{k})\}$ converges to $\psi({\bm{x}}^{*})$, for some ${\bm{x}}^{*}\in\mathcal{A}$.
Except for Lemma 2.16, similar results under slightly different conditions can be found in [47, 37, 48], while some recent works [60, 44] focus on analyzing the convergence of (2.3) under relaxed conditions. Interested readers could refer to those works for details.
In this section, we aim to establish the convergence properties of proposed framework (DSM) under two noise settings. To this end, it is crucial to ensure the uniform boundedness of iterates $\{{\bm{Z}}_{k}\}$. A large amount of literature [38, 41, 37, 42] regards this condition as a prior assumption when analyzing the convergence of SGD-type algorithms in a single-agent setting. Nevertheless, more recent works [61, 49, 62, 63, 64, 43] focus on rigorously establishing the uniform boundedness of the iterative sequence.
Section 3.1 introduces the fundamental definitions and assumptions used in our analysis, and then rigorously demonstrate the uniform boundedness and asymptotic convergence for our framework (DSM). In Sections 3.2 and 3.3, we further discuss about the convergence properties of (DSM) when the evaluation noise is introduced by random reshuffling and with-replacement sampling, respectively.
3.1 Basic assumptions and main results
(1.5) shows that the average update direction $\{\frac{1}{d}{\bm{H}}_{k}{\bm{1}_{d}}\}$ approximates $\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}({\bm{z}}_{i,k})$. To constrain the behavior of $\{\Phi_{i}\}_{i=1}^{d}$, we impose a set of basic assumptions concerning both the framework in Section 1.3 and its associated continuous-time differential inclusion,
$\frac{\mathrm{d}{\bm{z}}}{\mathrm{d}t}\in-\Phi({\bm{z}}):=\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}({\bm{z}})\right).$ (3.1)